The vapor pressure of pure water at 26◦C is 25.21 torr. What is the vapor pressure of
a solution that contains 20g glucose C6H12O6 in 70g of water?
1
Expert's answer
2020-11-20T09:59:48-0500

RMM of C6H12O6= 180g/mol

RMM of H2O=18g/mol

Moles of glucose= 20.0/180

=0.1111moles

moles of H2O=70/18

=3.89moles

Mole fraction of glucose= 0.1111/(0.1111+3.89)=0.0278

(P0-P)/P0=X, where X= 0.0278

P=25.21-(25.21*0.0278)= 24.5torr